# Curvature Tuning: Provable Training-free Model Steering From a Single Parameter

### v4

@@ C.1 CT Operates as a Projection
At its core, Curvature Tuning operates by modulating the non-linearity of the activation functions of a trained model, providing a novel approach to model steering. In order to formalize the effect of CT, the following briefly introduces the notion of spaces of smooth functions.
@@
[…]
@@
We begin by considering the Sobolev semi-norm of a ReLU network (equivalent to the case of Eq. 8 with $\beta\to 1$). For each $\mathbf{x}\in\mathbb{R}^{d}$, the gradient of a ReLU network
@@
with $\varphi(z):=\max(0,z)$, for $z\in\mathbb{R}$, is given by
@@
where $D^{\ell}(\mathbf{x})$ is a diagonal matrix with $D^{\ell}_{ii}(\mathbf{x})=\mathbf{1}_{\{\mathbf{z}^{\ell}_{i}>0\}}$, with $\mathbf{z}_{i}^{\ell}=W^{\ell}_{i}\mathbf{z}^{\ell-1}+\mathbf{b}^{\ell}_{i}$ denoting the pre-activation of the $\ell$-th layer, for $\ell=1,\ldots,L$, with $\mathbf{z}^{0}:=\mathbf{x}$.
@@
Equipped with the above observations, in the following we characterize CT by formally restating and proving Theorem 3.1.
@@
Particularly, it holds $\|\nabla_{\mathbf{x}}^{2}f(\mathbf{x})\|_{L^{2}(\Omega)}\leq\|\nabla^{2}_{\mathbf{x}}f_{\beta,c}(\mathbf{x})\|_{L^{2}(\Omega)}$, from which $f_{\beta,a}$ enjoys higher local expressivity (non-vanishing curvature), while retaining the same model parameter $\mathbf{W}$.
Before proving Theorem C.1, we state the following Lemma, bounding the derivative of a CTU.
@@
Let $\varphi_{\beta,c}(x)$ be defined according to Eq. 8, for $\beta\in{[}0,1)$ and $c\in{[}0,1{]}$. Then
@@
$-c\overline{h}_{b}\leq\varphi_{\beta,c}^{\prime}(x)\leq 1+c\overline{h}_{b}\qquad\forall x\in\mathbb{R},\quad\beta\in{[}0,1)$ (18)
@@
For the case $x<0$, by using the identity $\sigma(x)=1-\sigma(-x)$, we have that $-\overline{h}_{b}\leq\varphi_{\beta,1}^{\prime}(x)\leq 1$. By combining the results, we have
$-\overline{h}_{b}\leq\varphi_{\beta,1}^{\prime}(x)\leq 1+\overline{h}_{b}\qquad\forall x\in\mathbb{R},\quad\beta\in{[}0,1)$ (19)
In conclusion, by convex combination of cases $c=0$ and $c=1$, Eq. 19 holds uniformly in $x$. ∎
We can now prove Theorem C.1. To do so, for $f_{\beta,c}$ we have to show that
@@
where $D^{\ell}_{\beta,c}(\mathbf{z^{\ell}})=\operatorname{diag}(\varphi_{\beta,c}^{\prime}(\mathbf{z}^{\ell}))$ with $\varphi_{\beta,c}^{\prime}(\mathbf{z}^{\ell})_{i}:=\varphi_{\beta,c}^{\prime}(\mathbf{z}^{\ell}_{i})$ according to Eq. 17.
@@
Importantly, since a ReLU network $f$ has vanishing curvature a.e., then for $10\leq\beta<$ 1, we have
@@
Theorem C.1 shows that CT operates by projecting a ReLU network $f$ to a smooth function $f_{\beta,c}$ in a restricted Sobolev space. Crucially, $f_{\beta,c}$ enjoys bounded gradients (and so is well behaved), and non-vanishing local-curvature for $10<\beta<$ 1, making it locally more expressive than the affine spline $f$, for fixed $\mathbf{W}$.
@@
We leave for future work extending our result to Train CT, which is associated with a non-convex optimization problem of finding optimal $(\beta,c)$ for every neuron in the network. An additional important direction is to more closely compare $\|\nabla_{\mathbf{x}}f\|$ and $\|\nabla_{\mathbf{x}}f_{\beta,c}\|$, which may reveal more precise Lipschitz behaviour for CT, potentially better guiding the search for $\beta$ and $c$.
@@ C.2 Toy Example
We conclude the discussion by providing the full derivation for the motivating example in Section 3.
@@
[…]
with $z:=\mathbf{1}_{\{\mathbf{A}_{k,\cdot}\bm{\gamma}(t)+\mathbf{b}_{k}<0\}}$. Then,
[…]
$=2\pi\sum\limits_{k=1}^{r^{\prime}-1}\int_{t_{k}}^{t_{k+1}}(-1)^{z}\left(\mathbf{A}_{k,1}\frac{\sin 2\pi}{2\pi}-\mathbf{A}_{k,2}\frac{\cos 2\pi}{2\pi}+\mathbf{b}_{k}t\right)_{t_{k}}^{t_{k+1}}$ (29)
which evaluates to
$e$ $=\sum\limits_{k=1}^{r^{\prime}-1}(-1)^{z}\Bigg{(}2\pi\mathbf{b}_{k}(t_{k+1}-t_{k})+$ (31)
$+\mathbf{A}_{k,1}\left(2\sin{\frac{t_{k+1}-t_{k}}{2}}\cos{\frac{t_{k+1}-t_{k}}{2}}\right)-\mathbf{A}_{k,2}\left(2\sin{\frac{t_{k+1}+t_{k}}{2}}\sin{\frac{t_{k}-t_{k+1}}{2}}\right)\Bigg{)}$
from which clearly $e\to 0\iff t_{k+1}\to t_{k}\quad\forall k$.
Hence, assuming the ReLU network considered attained optimal approximation error $e>0$, reducing the error further requires increasing the number of breakpoints of the ASO, in turn requiring a degree of retraining (either through PEFT or training from scratch). With this view, Curvature Tuning opens an additional avenue for model adaptation: steering the model’s decision boundaries by modulating the non-linearity of the activation function, allowing to tune a model towards optimality without expensive retraining. To this end, it is important to note that modulating decision boundaries is orthogonal to feature adaptation and finetuning, since it allows to change the shape of decision boundaries while keeping the model parameter $\mathbf{W}$ fixed.
@@ Appendix D Curvature Tuning (CT) implementation
The following code provides the Python implementation for CT and Trainable CT:
  • CTU & TrainableCTU: classes that define the CTU module used in CT and Trainable CT, respectively.
  • replace_module & replace_module_dynamic: functions that apply the appropriate module replacement to integrate CT or Trainable CT into a model.
@@
from torch import nn
@@
class CTU(nn.Module):
CTU for CT.
@@
class TrainableCTU(nn.Module):
@@
self._raw_beta = nn.Parameter(torch.full(param_shape, float(raw_beta)))
@@
self._raw_coeff = nn.Parameter(torch.full(param_shape, float(raw_coeff)))
@@
def replace_module(model, old_module=nn.ReLU, new_module=CTU, **kwargs):
@@
def replace_module_dynamic(model, input_shape, old_module=nn.ReLU, new_module=TrainableCTU, **kwargs):
@@ Appendix E LoRA Implementation
The following code provides the Python implementation of LoRA used in Section 4:
@@
from torch import nn as nn
@@
class LoRALinear(nn.Module):
@@
def __init__(self, original_layer: nn.Linear, r: int = 4, alpha: float = 1.0):
@@
self.weight = nn.Parameter(original_layer.weight.data, requires_grad=False)
@@
self.bias = nn.Parameter(original_layer.bias.data, requires_grad=False)
@@
self.B = nn.Parameter(torch.zeros((self.out_features, r)))
self.A = nn.Parameter(torch.zeros((r, self.in_features)))
@@
nn.init.kaiming_uniform_(self.B, a=5 ** 0.5)
nn.init.zeros_(self.A)
@@
class LoRAConv2d(nn.Module):
@@
def __init__(self, original_layer: nn.Conv2d, r: int = 4, alpha: float = 1.0):
@@
self.weight = nn.Parameter(original_layer.weight.data, requires_grad=False)
@@
self.bias = nn.Parameter(original_layer.bias.data, requires_grad=False)
@@
self.B = nn.Parameter(torch.zeros((self.out_channels, r)))
self.A = nn.Parameter(torch.zeros((r, fan_in)))
@@
nn.init.kaiming_uniform_(self.B, a=5 ** 0.5)
nn.init.zeros_(self.A)
@@
def get_lora_model(model: nn.Module, r: int = 4, alpha: float = 1.0):
@@
if isinstance(child, nn.Conv2d):
@@
elif isinstance(child, nn.Linear):

### v5

@@ C.1 CT operates as a projection
At its core, Curvature Tuning operates by modulating the nonlinearity of the activation functions of a trained model, providing a novel approach to model steering. In order to formalize the effect of CT, the following briefly introduces the notion of spaces of smooth functions.
@@
$\|f\|_{L^{p}(\Omega)}:=\left(\int_{\Omega}|f(\mathbf{x})|^{p}d\mathbf{x}\right)^{\frac{1}{p}}<\infty$ (10)
@@
We begin by considering the Sobolev semi-norm of a ReLU network (equivalent to the case of Eq.˜8 with $\beta\to 1$). For each $\mathbf{x}\in\mathbb{R}^{d}$, the gradient of a ReLU network
@@
with $\varphi(a)]_{i}:=\max(0,a_{i})$ for $a\in\mathbb{R}^{m}$ and $i\in[1,m]$, is given by
@@
where $D^{\ell}(\mathbf{x})$ is a diagonal matrix with $D^{\ell}_{ii}(\mathbf{x})=\mathbf{1}_{\{\mathbf{z}^{\ell}_{i}>0\}}$, with $\mathbf{z}_{i}^{\ell}=W^{\ell}_{i}\varphi(\mathbf{z}^{\ell-1})+\mathbf{b}^{\ell}_{i}$ denoting the pre-activation of the $\ell$-th layer, for $\ell=1,\ldots,L$, with $\mathbf{z}^{0}:=\mathbf{x}$.
@@
Equipped with the above observations, in the following we characterize CT by formally restating and proving Theorem˜3.1.
@@
Particularly, it holds $\|\nabla_{\mathbf{x}}^{2}f(\mathbf{x})\|_{L^{2}(\Omega)}\leq\|\nabla^{2}_{\mathbf{x}}f_{\beta,c}(\mathbf{x})\|_{L^{2}(\Omega)}$, from which $f_{\beta,c}$ enjoys higher local expressivity (non-vanishing curvature), while retaining the same model parameter $\mathbf{W}$.
Before proving Theorem˜C.1, we state the following Lemma, bounding the derivative of a CTU.
@@
Let $\varphi_{\beta,c}(x)$ be defined according to Eq.˜8, for $\beta\in{[}0,1)$ and $c\in{[}0,1{]}$. Then
@@
$c\overline{h}_{b}\leq\varphi_{\beta,c}^{\prime}(x)\leq 1+c\overline{h}_{b}\qquad\forall x\in\mathbb{R},\quad\beta\in{[}0,1)$ (18)
@@
For the case $x<0$, by using the identity $\sigma(x)=1-\sigma(-x)$, we have that $\overline{h}_{b}\leq\varphi_{\beta,1}^{\prime}(x)\leq 1$. By combining the results, we have
$\overline{h}_{b}\leq\varphi_{\beta,1}^{\prime}(x)\leq 1+\overline{h}_{b}\qquad\forall x\in\mathbb{R},\quad\beta\in{[}0,1)$ (19)
In conclusion, by convex combination of cases $c=0$ and $c=1$, Eq.˜19 holds uniformly in $x$. ∎
We can now prove Theorem˜C.1. To do so, for $f_{\beta,c}$ we have to show that
@@
where $D^{\ell}_{\beta,c}(\mathbf{z^{\ell}})=\operatorname{diag}(\varphi_{\beta,c}^{\prime}(\mathbf{z}^{\ell}))$ with $\varphi_{\beta,c}^{\prime}(\mathbf{z}^{\ell})_{i}:=\varphi_{\beta,c}^{\prime}(\mathbf{z}^{\ell}_{i})$ according to Eq.˜17.
@@
Importantly, since a ReLU network $f$ has vanishing curvature a.e., then for $0\leq\beta<1$, we have
@@
Theorem C.1 shows that CT operates by projecting a ReLU network $f$ to a smooth function $f_{\beta,c}$ in a restricted Sobolev space. Crucially, $f_{\beta,c}$ enjoys bounded gradients (and so is well behaved), and non-vanishing local-curvature for $0<\beta<1$, making it locally more expressive than the affine spline $f$, for fixed $\mathbf{W}$.
@@
We leave for future work extending our result to T-CT, which is associated with a non-convex optimization problem of finding optimal $(\beta,c)$ for every neuron in the network. An additional important direction is to more closely compare $\|\nabla_{\mathbf{x}}f\|$ and $\|\nabla_{\mathbf{x}}f_{\beta,c}\|$, which may reveal more precise Lipschitz behaviour for CT, potentially better guiding the search for $\beta$ and $c$.
@@ C.2 Toy example
We conclude the discussion by providing the full derivation for the motivating example in Section˜3.
@@
Then, the error is expressed by the line integral $e=\int_{\bm{\gamma}}|f|d\mathbf{x}=\int_{0}^{1}|f(\bm{\gamma}(t))|\|\bm{\gamma}^{\prime}(t)\|dt$. Since $f$ is an Affine Spline Operator, and each linear region in $\Omega$ is convex, then the integral along $\bm{\gamma}$ can be broken down into the integral along the intersection of $\bm{\gamma}$ with the spline partition $\Omega$, i.e. $\Omega_{\bm{\gamma}}:=\Omega\cap S^{1}$. Importantly, this allows us to pull back the affine spline breakpoints from $\Omega_{\bm{\gamma}}$ to [0,1]{[}0,1{]}, so that $0\leq t_{1}\leq\ldots\leq t_{r^{\prime}}\leq 1$, where $r^{\prime}=|\Omega_{\bm{\gamma}}|$. And we augment the breakpoints with the end points so that $0=t_{0}\leq t_{1}\leq\ldots\leq t_{r^{\prime}}\leq t_{r^{\prime}+1}=1$. Then,
$e$ $=\int_{0}^{1}|f(\bm{\gamma}(t))|\|\bm{\gamma}^{\prime}(t)\|dt$ (25)
$=2\pi\sum\limits_{k=0}^{r^{\prime}}\int_{t_{k}}^{t_{k+1}}|\mathbf{A}_{r_{k},\cdot}\bm{\gamma}(t)+\mathbf{b}_{r_{k}}|dt$ (26)
$=2\pi\sum\limits_{k=0}^{r^{\prime}}\int_{t_{k}}^{t_{k+1}}(-1)^{z_{k}(t)}\left(\mathbf{A}_{r_{k},\cdot}\bm{\gamma}(t)+\mathbf{b}_{r_{k}}\right)dt$ (27)
with $z_{k}(t):=\mathbf{1}_{\{\mathbf{A}_{r_{k},\cdot}\bm{\gamma}(t)+\mathbf{b}_{r_{k}}<0\}}$, where $r_{k}$ denotes which spline region the $k$-th segment $t_{k},t_{k+1}]$ falls into. Then,
$e$ $=2\pi\sum\limits_{k=0}^{r^{\prime}}\int_{t_{k}}^{t_{k+1}}(-1)^{z_{k}(t)}\left(\mathbf{A}_{r_{k},1}\cos 2\pi t+\mathbf{A}_{r_{k},2}\sin 2\pi t+\mathbf{b}_{r_{k}}\right)dt$ (28)
$=2\pi\sum\limits_{k=0}^{r^{\prime}}\left(\int_{t_{k}}^{s_{k}}(-1)^{z_{k}(t)}g^{\prime}_{r_{k}}(t)dt+\int_{s_{k}}^{t_{k+1}}(-1)^{z_{k}(t)}g^{\prime}_{r_{k}}(t)dt\right)$ (29)
$=2\pi\sum\limits_{k=0}^{r^{\prime}}\left((-1)^{z_{k}(t_{k})}\left[g_{r_{k}}(t)\right]_{t_{k}}^{s_{k}}+(-1)^{z_{k}(t_{k+1})}\left[g_{r_{k}}(t)\right]_{s_{k}}^{t_{k+1}}\right)$ (30)
where
$g^{\prime}_{r_{k}}(t)=\mathbf{A}_{r_{k},1}\cos 2\pi t+\mathbf{A}_{r_{k},2}\sin 2\pi t+\mathbf{b}_{r_{k}},$
$g_{r_{k}}(t)=\mathbf{A}_{r_{k},1}\frac{\sin 2\pi t}{2\pi}-\mathbf{A}_{r_{k},2}\frac{\cos 2\pi t}{2\pi}+\mathbf{b}_{r_{k}}t,$
and $s_{k}\in[t_{k},t_{k+1}]$ is defined so $z_{k}(t)$ holds the same value for $t\in[t_{k},s_{k}]$ and the opposite for $t\in(s_{k},t_{k+1}]$. If for $t\in[t_{k},t_{k+1}]$, $z_{k}(t)$ holds the same value, then simply set $s_{k}=t_{k}$.
Then since both $(-1)^{z_{k}(t_{k})}\left[g_{r_{k}}(t)\right]_{t_{k}}^{s_{k}}$ and $(-1)^{z_{k}(t_{k+1})}\left[g_{r_{k}}(t)\right]_{s_{k}}^{t_{k+1}}$ are non-negative, it is clear $t_{k+1}\to t_{k}\quad\forall k\implies e\to 0$.
Hence, assuming the ReLU network considered attained optimal approximation error $e>0$, reducing the error further requires increasing the number of breakpoints of the ASO, in turn requiring a degree of retraining (either through PEFT or training from scratch). With this view, Curvature Tuning opens an additional avenue for model adaptation: steering the model’s decision boundaries by modulating the nonlinearity of the activation function, allowing to tune a model towards optimality without expensive retraining. To this end, it is important to note that modulating decision boundaries is orthogonal to feature adaptation and finetuning, since it allows to change the shape of decision boundaries while keeping the model parameter $\mathbf{W}$ fixed.
@@ Appendix D Curvature Tuning (CT) implementation
The following code provides the Python implementation for S-CT and T-CT:
  • SCTU & TCTU: classes that define the CTU module used in S-CT and T-CT, respectively.
  • replace_module & replace_module_dynamic: functions that apply the appropriate module replacement to integrate S-CT or T-CT into a model.
@@
from torch import $n$
@@
class SCTU($n$.Module):
CTU for Steering CT.
@@
class TCTU($n$.Module):
@@
self._raw_beta = $n$.Parameter(torch.full(param_shape, float(raw_beta)))
@@
self._raw_coeff = $n$.Parameter(torch.full(param_shape, float(raw_coeff)))
@@
def replace_module(model, old_module=$n$.ReLU, new_module=SCTU, **kwargs):
@@
def replace_module_dynamic(model, input_shape, old_module=$n$.ReLU, new_module=TCTU, **kwargs):
@@ Appendix E LoRA implementation
The following code provides the Python implementation of LoRA used in Section˜4:
@@
from torch import $n$ as $n$
@@
class LoRALinear($n$.Module):
@@
def __init__(self, original_layer: $n$.Linear, r: int = 4, alpha: float = 1.0):
@@
self.weight = $n$.Parameter(original_layer.weight.data, requires_grad=False)
@@
self.bias = $n$.Parameter(original_layer.bias.data, requires_grad=False)
@@
self.B = $n$.Parameter(torch.zeros((self.out_features, r)))
self.A = $n$.Parameter(torch.zeros((r, self.in_features)))
@@
$n$.init.kaiming_uniform_(self.B, a=5 ** 0.5)
$n$.init.zeros_(self.A)
@@
class LoRAConv2d($n$.Module):
@@
def __init__(self, original_layer: $n$.Conv2d, r: int = 4, alpha: float = 1.0):
@@
self.weight = $n$.Parameter(original_layer.weight.data, requires_grad=False)
@@
self.bias = $n$.Parameter(original_layer.bias.data, requires_grad=False)
@@
self.B = $n$.Parameter(torch.zeros((self.out_channels, r)))
self.A = $n$.Parameter(torch.zeros((r, fan_in)))
@@
$n$.init.kaiming_uniform_(self.B, a=5 ** 0.5)
$n$.init.zeros_(self.A)
@@
def get_lora_model(model: $n$.Module, r: int = 4, alpha: float = 1.0):
@@
if isinstance(child, $n$.Conv2d):
@@
elif isinstance(child, $n$.Linear):